OpenStudy (anonymous):

What is the slope of the line that contains the points (-2, 1) and (0, -3)?

OpenStudy (anonymous):

do you know the formula of slope of a line passing through (a.b) and(c,d)??

OpenStudy (anonymous):

nope i dont

OpenStudy (anonymous):

the formula is (difference of y coordinates)/difference of x coordinates,i.e, (d -b)/(c-a) Just plug in the values. The answer is -2

OpenStudy (anonymous):

\[(y2-y1)/(x2-x1)\] (x1, y1), (x2, y2)=(-2, 1), (0, -3) x1=-2, y1=1, x2=0, y2=-3 Like @Rohith15 said, just plug in the values and you should get -2 :3 Let me know if that makes sense.

OpenStudy (anonymous):

@girlnotonfire the slope is -2 right??

OpenStudy (anonymous):

Yes! (-3-1)/(0-(-2)) =-4/2 =-2 :)